In 2018, the net income was $3600 and cash dividends was $824. In 2019, the net income was $6758 and cash dividends was $1260.

1. What is the dividend payout ratio in 2018 and 2019?

2. What is the plowback ratio in 2018 and 2019?
Answer #1

Dividend payout ratio=Dividend payout/Net income

Plowback ratio-1-Dividend payout ratio

2018 2019
Dividend payout ratio (824/3600)=22.89%(or 0.2289)(Approx) (1260/6758)=18.64%(or 0.1864 approx)
Plowback ratio (1-0.2289)=77.11%(or 0.7711 approx) (1-0.1864)=81.36%(or 0.8136 approx).